# Leveraging the true depth of LLMs

### v1

@@ Abstract
Large Language Models demonstrate remarkable capabilities at the cost of high compute requirements. While recent research has shown that intermediate layers can be removed or have their order shuffled without impacting performance significantly, these findings have not been employed to reduce the computational cost of inference. We investigate several potential ways to reduce the depth of pre-trained LLMs without significantly affecting performance. Leveraging our insights, we present a novel approach that exploits this decoupling between layers by grouping some of them into pairs that can be evaluated in parallel. This modification of the computational graph—through better parallelism—results in an average improvement of around 1.20x on the number of tokens generated per second, without re-training nor fine-tuning, while retaining 95%-99% of the original accuracy. Empirical evaluation demonstrates that this approach significantly improves serving efficiency while maintaining model performance, offering a practical improvement for large-scale LLM deployment.
@@ 1 Introduction
Figure 1: Trade-off between perplexity and inference time of Llama 2 7B and Llama 3.2 3B with our novel Layer Parallelism. $\Delta$ indicates the number of consecutive layers that we parallelize in pairs. The average time is the wall clock time to generate 4096 tokens (with KV-Cache) on x2 A100 SXM4 80Gb. The perplexity is measured on RedPajama (Together Computer, 2023). As shown by the gap between the two Pareto fronts (black arrow), applying our approach to Llama 2 7b results in both faster generation speeds and better performance than the vanilla Llama 3.2 3B.
@@
The rapid advancement of Large Language Models (LLMs) and deep neural networks has revolutionized Artificial Intelligence applications across industries. However, the ever-increasing computational demands of these models, with parameters often numbering in the hundreds of billions, present significant commercial challenges. Efficient inference is crucial for organizations deploying these models at scale, as it directly impacts operational costs, user experience, and environmental sustainability (Singh et al., 2025; Xu et al., 2024; Wu et al., 2022). Monthly cloud computing expenses for LLM inference can reach millions of dollars for high-traffic applications, making optimization techniques essential. Additionally, reducing inference latency is critical for real-time applications and for deploying models on devices with more limited compute. Thus, developing and implementing efficient inference methods has become a key differentiator in the competitive AI marketplace, driving both innovation and profitability.
@@
Large Language Models (LLMs) have evolved to incorporate architectures with hundreds of layers (OpenAI, 2023; et. al., 2024). These models are constructed from stacked transformer blocks, each comprising attention and feed-forward sub-blocks, with a residual stream traversing the entire architecture to facilitate efficient gradient backpropagation during training. This architectural choice parallels the design principles of ResNets (He et al., 2015), where research has demonstrated that network depth may be partially redundant, allowing for layer reordering without significant performance degradation (Veit et al., 2016). Recent investigations have revealed similar flexibility in transformer architectures (Lad et al., 2024), where interventions such as layer deletion and swapping are applied without large performance degradations. While these findings challenge our understanding of LLMs’ effective depth requirements, their potential for optimizing inference efficiency remains unexplored.
@@
Our research leverages these insights to investigate various approaches for reducing the depth of LLMs. We explore several intervention strategies to modify the computational graph of pre-trained LLMs, such as shuffling, pruning or merging layers. A key observation is that the demonstrated ability to reorder blocks enables parallel execution strategies. Specifically, we demonstrate that multiple consecutive block pairs can be processed in parallel while maintaining acceptable performance metrics across perplexity and In-Context Learning (ICL) benchmarks. We introduce a novel approach for layer parallelization that enhances inference speed, with a minimal drop in performance. Furthermore, we show that the performance degradation observed across various benchmarks can be partially mitigated through targeted fine-tuning procedures.
@@
  • We find that we can define a parallelization transform on the computational graph of two sequential Transformer layers, and stack this parallelization operation to several sequential pairs of layers without loosing significant performance. Our approach can be applied to existing Transformer models and does not require re-training.
  • We exploit this parallelization of the computational graph to run the models around 1.20x faster using multiple GPUs without loosing much performance and ICL capabilities.
  • We show that by fine-tuning a parallelized model we can recover some of the lost performance, while retaining the previously obtained speed-up.
The source code can be found here: https://github.com/noctrog/effective-depth.
+FFNLNMHALN++FFN1LNFFN2LNMHA1LNMHA2LN+(a)(b)
Figure 2: Comparing a normal transformer block with our layer parallel implementation. In (a) we show a normal Transformer Layer. In (b) we illustrate our layer parallelization, where each column runs separately on one or multiple GPUs. The residuals are gathered and synchronized through a reduce operation across all GPUs. The pre-attention LayerNorms’ weights are copied from the original blocks, while the post-attention LayerNorm’s weights are averaged and shared in both layers.
@@ 2 Related work
The effective depth of Deep Networks. Theoretically, any feed-forward network with at least one hidden layer can model any function, given enough width (Pinkus, 1999). In practice, it is easier to achieve high expressivity by increasing the model’s depth. But naively increasing the depth can make things more difficult for the optimizer, since the gradients now have to flow through many layers. To alleviate this problem, ResNets (He et al., 2015) introduced skip connections at regular intervals to allow an easy flow of the gradient to the first layers. Alternatively, in Inception (Szegedy et al., 2014), the researchers investigated ways to boost computational power by adding additional processing units along different parallel pathways in the computational network, rather than just along a single sequential path. A unification of both methods can be found in the Highway Networks (Srivastava et al., 2015), where the skip connection of the residual blocks consists of another block of compute. Nowadays, residual connections are ubiquitous in large models.
@@
Figure 3: Changes in perplexity when applying transformations on contiguous stretches of layers. Each one of the five heatmaps above correspond to a transformation of a group of consecutive layer, where the row index $s$ corresponds to the first layer of the group, and the column index $e$ to the last. The color coding indicates how the perplexity—estimated on a subset of RedPajama (Together Computer, 2023)—is impacted by the corresponding modification of the model. The perplexity for the base Llama 2 7B model is $6.2$. In (a), we shuffle—for each forward—the layers from $s$ to $e$. We can see that many consecutive layers can be shuffled with little impact on the overall perplexity. For instance, shuffling layers $15$ to 2525252510101010 layers in total—raises the perplexity only to $9.1$. In (b), we prune contiguous stretches of layers. We can see that not many blocks can be removed without starting to significantly degrade the perplexity. In (c) we merge contiguous layers. The results with merging are near identical to those for pruning. This reveals there is no advantage in merging layers, most likely a results of averaging matrices not originating from the same initial values. In (d) we run contiguous blocks in parallel. Given the success of shuffling, it makes sense that this approach works well. Running blocks $17$ to $27$ raises the perplexity to $9.3$. Finally, in (e) we run pairs of consecutive layers in parallel. As a result, we can parallelize much longer stretches of layers. For instance, we can apply this transformation from layer $4$ to $29$ and only increase the perplexity to $9.1$. This reduces the depth of the model from $32$ to $19$. This result makes it possible for us to leverage this parallelism for faster inference as we discuss in § 4.
@@
Efficient inference of LLMs. Several complementary approaches exist for enhancing the computational efficiency of large-scale models, primarily through pruning and sparsity, quantization, and parallelism. Pruning(LeCun et al., 1989; Han et al., 2015, 2016; Frantar & Alistarh, 2023) constitutes a dimensional reduction methodology that systematically eliminates redundant parameters while preserving model performance, thereby introducing architectural sparsity. This methodology is founded on empirical evidence demonstrating that neural networks frequently exhibit overparameterization, containing numerous weights with negligible contribution to the output. Through sophisticated pruning strategies, the inherent sparsity support in contemporary accelerators can be leveraged to enhance both memory utilization and computational efficiency (Zhang et al., 2020; Wang et al., 2021). In contrast, quantization encompasses the transformation of floating-point numerical representations (predominantly FP32) into reduced-precision integer formats, such as INT8 or INT4 (Han et al., 2016; Jacob et al., 2018). When implemented on hardware accelerators, these lower-precision representations facilitate superior memory bandwidth utilization, addressing a primary bottleneck in modern large-scale models (Gholami et al., 2024); moreover, integer-based computations yield enhanced processing speed and substantially improved energy efficiency (Horowitz, 2014). Finally, parallelization techniques during inference, such as tensor and pipeline parallelism, enable the distribution of computational workload across multiple devices, thereby reducing latency and increasing throughput, although this often requires careful consideration of communication overhead and load balancing (Li et al., 2024; Narayanan et al., 2021).
@@
Parallelism via Computational Graph Optimization. Recent research has investigated architectural layer-level optimization strategies to enhance transformer model inference efficiency. The Staircase Transformer (Cutler et al., 2025) implements parallel layer execution with dynamic recurrent computation based on model requirements. Similarly, the Staggering Transformer (Cai et al., 2024) achieves layer parallelization by connecting layer $l_{k}$ at time step $t$ to both the $(t-1)$ output of layer $l_{k-1}$ and the $t$ output of layer $l_{k-2}$. To the best of our knowledge, no research has addressed the fusion of consecutive layers through tensor parallelism.
@@ 3 Effective Depth
In this section, we investigate the effective depth of LLMs. By applying several transformations to a pretrained Transformer LLM, and measuring the resulting perplexity degradation, we reveal the loose dependencies between intermediary layers. The transformations consist of shuffling, merging, and pruning transformer layers. To avoid the combinatorial explosion resulting from considering all the possible subsets of transformer layers, we instead apply our transformations to all the contiguous stretch of layers. If $L=\{\ell_{1},\ldots,\ell_{N}\}$ are the ordered layers, then we apply our transformations to all the sub-list $\{\ell_{i}\}_{i=s}^{e}$ with $1\leq s\leq e\leq N$. Previous works have shown that—at least when considering pruning—the importance of layers is well behaved, with low importance layers close to one another (Men et al., 2024), which justifies considering contiguous stretch of layers only.
Shuffling blocks. We start by shuffling contiguous stretches of layers, re-ordering the layers according to random permutations. Results are shown in Fig. 3.(a). While shuffling the early and two last layers significantly raises the perplexity, there are large stretches of blocks which can be shuffled with surprisingly low impact on the perplexity. For instance, one can shuffle the layers $\{\ell_{i}\}_{15\leq i<25}$ and only have a increase in perplexity of $2.9$. This goes against the classical belief that models build deeply hierarchical representations, where features in previous layers are leveraged to build more complex features in later layers. We interpret this as multiple layers working at the same level of abstraction. Using these insights, we define the effective depth of an LLM as the shortest depth required to efficiently leverage existing latent representations without a significant loss of performance. This reveals an important level of layer decoupling within the model.
Running blocks in parallel. The observed layer decoupling suggests that specific transformer operations may be executed independently, providing an opportunity for parallel computation. More precisely, let’s consider two sequential transformer layers $\ell_{k}$ and $\ell_{k+1}$, each comprising attention and FFN sub-blocks ($A_{k}(\cdot)$ and $F_{k}(\cdot)$, respectively). The standard sequential output ${\bm{y}}$ for these layers, given an input ${\bm{x}}$, is given by:
@@
Contiguous 2-parallel. Instead of parallelizing long stretches of layers, we experiment with running pairs of consecutive layers in parallel. This springs from the assumption that local ordering matters less than global ordering, i.e. shuffling consecutive layers is less risky than shuffling layers further apart. As an example, if we apply the proposed transformation to layers $\{\ell_{15},\ell_{16},\ell_{17},\ell_{18},\ell_{19}\}$, it would result in the following process: (1) the two layers $\{\ell_{15},\ell_{16}\}$ process the input in parallel (according to equation (PAR)), (2) the output is forwarded to layers $\{\ell_{17},\ell_{18}\}$ which process it in parallel; finally, in (3) their joint output is fed to layer $\ell_{19}$ which processes it on its own as any normal layer. The effect of such transformation of the compute graph on the perplexity can be seen in Fig. 3(e). Remarkably, it is possible to run wide stretches of consecutive pairs of blocks in parallel with only a minor degradation of perplexity. For instance, one can apply this transformation from layer $4$ to layer $29$ with only a degradation of perplexity of $2.9$, while reducing the model depth from $32$ to $19$. The success of this approach led us to also try running triplets of consecutive layers in parallel, but we found it to perform less well.
Exploring other transformations. We also experiment with pruning (Fig. 3(b)) and merging (Fig. 3(c)). Pruning has already been studied in several prior works (Gromov et al., 2024; Jung et al., 2019). While pruning a layer leads to a larger perplexity increase compared to parallelizing two blocks, it offers the advantage of reducing the model’s parameter count, which results in improved throughput and more efficient memory usage.

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities at the cost of high compute requirements. Recent studies have demonstrated that intermediate layers in LLMs can be removed or reordered without substantial accuracy loss; however, this insight has not yet been exploited to improve inference efficiency. Leveraging observed layer independence, we propose a novel method that groups consecutive layers into pairs evaluated in parallel, effectively restructuring the computational graph to enhance parallelism. Without requiring retraining or fine-tuning, this approach achieves an inference throughput improvement of 1.05x–1.20x on standard benchmarks, retaining 95%–99% of the original model accuracy. Empirical results demonstrate the practicality of this method in significantly reducing inference cost for large-scale LLM deployment. Additionally, we demonstrate that modest performance degradation can be substantially mitigated through lightweight fine-tuning, further enhancing the method’s applicability.
@@ 1 Introduction
Figure 1: The effect of LP on execution time (4K tokens) and perplexity (measured against RedPajama[1]). The bigger Llama model achieves both better perplexity and faster execution speeds than the smaller Llama model.
@@
The rapid advancement of LLMs has revolutionized Artificial Intelligence applications across industries. However, the ever-increasing computational demands of these models, with parameters often numbering hundreds of billions, present significant commercial challenges. Efficient inference is crucial for organizations that deploy these models at scale, as it directly impacts operational costs, user experience, and environmental sustainability [2, 3, 4]. Monthly cloud computing expenses for LLM inference can reach millions of dollars for high-traffic applications, making optimization techniques essential. In addition, reducing inference latency is critical for real-time applications and for deploying models on devices with more limited compute. Thus, the development and implementation of efficient inference methods has become a key differentiator in the competitive AI market, driving both innovation and profitability.
+FFNLNMHALN++FFN1LNFFN2LNMHA1LNMHA2LN+(a)(b)
Figure 2: Comparing a normal transformer block with our layer parallel implementation. In (a) we show a normal Transformer Layer. In (b) we illustrate our layer parallelization, where each column runs separately on one or multiple GPUs. The residuals are gathered and synchronized through a reduce operation across all GPUs. The LayerNorms’ weights are copied from the original blocks, and each divergent path uses the normalization from its original layer.
@@
LLMs have evolved to incorporate architectures with hundreds of layers [5, 6]. These models are constructed from stacked transformer blocks, each comprising attention and feedforward subblocks, with a residual stream traversing the entire architecture to facilitate efficient gradient backpropagation during training. This architectural choice parallels the design principles of ResNets [7], where research has shown that network depth may be partially redundant, allowing layer reordering without significant performance degradation [8]. Recent investigations have revealed similar flexibility in transformer architectures [9], where interventions such as layer removal and swapping are applied without large performance degradations. Although these findings challenge our understanding of LLMs’ true effective depth, their potential for optimizing inference efficiency remains unexplored.
@@
Stemming from this line of work, we set to find approaches that reduce the depth of LLMs. We do so by exploring several interventions to the computational graph of pre-trained LLMs such as shuffling, pruning or merging layers. As a result of this search, we find that multiple consecutive block pairs can be processed in parallel while maintaining acceptable performance across perplexity and In-Context Learning (ICL) benchmarks. We introduce a novel parallelism approach, Layer Parallelism (LP), that enhances inference speed when making use of tensor parallelism with a minimal drop in performance. Furthermore, we show that the performance degradation observed across various benchmarks can be partially mitigated through targeted fine-tuning procedures.
@@
  • We find that we can define a parallelization transform on the computational graph of two sequential Transformer layers, and stack this parallelization operation to several sequential pairs of layers without loosing significant ICL performance. Our approach, which we call Layer Parallelism (LP), can be applied to existing Transformer models and does not require retraining.
  • We show that by fine-tuning the LP blocks we can recover some of the lost performance, while retaining the previously obtained speed-up.
As a result, the rearrangement of the computational graph allows for faster inference speeds when serving the LLM using tensor parallelism. The speed-up mainly arises from the decrease in the number of all-reduce operations between GPUs. This is because LP uses two all-reduce operations per pair of transformer decoder layers, while normal layers would require four all-reduce operations.
@@ 2 Related work
The effective depth of Deep Networks. Theoretically, given enough width, any feed-forward network with at least one hidden layer can model any function [10]. In practice, it is easier to achieve high expressivity by increasing the model’s depth. However, naively increasing network depth can complicate optimization, since the gradients now have to flow through many layers. To alleviate this problem, ResNets [7] introduced skip connections at regular intervals to allow an easy flow of the gradient to the first layers. Alternatively, Inception [11] explored approaches to boost computational power by adding additional processing units along different parallel pathways in the computational network, rather than just along a single sequential path. A unification of both methods can be found in the Highway Networks [12], where the skip connection of the residual blocks consists of another block of compute. Nowadays, residual connections are ubiquitous in large models.
@@
Figure 3: Changes in perplexity when applying transformations on contiguous stretches of layers. Each one of the five heatmaps above correspond to a transformation of a group of consecutive layer, where the row index $s$ corresponds to the first layer of the group, and the column index $e$ to the last. The color coding indicates how the perplexity—estimated on a subset of RedPajama [1]—is impacted by the corresponding modification of the model. The perplexity for the base Llama 2 7B model is $6.2$. In (a), we shuffle—for each forward—the layers from $s$ to $e$. We can see that many consecutive layers can be shuffled with little impact on the overall perplexity. For instance, shuffling layers $15$ to 2525252510101010 layers in total—raises the perplexity only to $9.1$. In (b), we prune contiguous stretches of layers. We can see that not many blocks can be removed without starting to significantly degrade the perplexity. In (c) we merge contiguous layers. The results with merging are near identical to those for pruning. This reveals there is no advantage in merging layers, most likely a results of averaging matrices not originating from the same initial values. In (d) we run contiguous blocks in parallel. Given the success of shuffling, it makes sense that this approach works well. Running blocks $17$ to $27$ raises the perplexity to $9.3$. Finally, in (e) we run pairs of consecutive layers in parallel. As a result, we can parallelize much longer stretches of layers. For instance, we can apply this transformation from layer $4$ to $29$ and only increase the perplexity to $9.1$. This reduces the depth of the model from $32$ to $19$. This result makes it possible for us to leverage this parallelism for faster inference as we discuss in § 4.
@@
Efficient inference of LLMs. Several complementary approaches exist for enhancing the computational efficiency of large-scale models, primarily through pruning/sparsity, quantization, and parallelism. Pruning[13, 14, 15, 16] constitutes a dimensional reduction methodology that systematically eliminates redundant parameters while preserving model performance, thereby introducing architectural sparsity. This methodology is founded on empirical evidence demonstrating that neural networks frequently exhibit overparameterization, containing numerous weights with negligible contributions to the output. Through sophisticated pruning strategies, the inherent sparsity support in contemporary accelerators can be leveraged to enhance both memory utilization and computational efficiency [17, 18]. Early-exit[19, 20] can be seen as a way of runtime layer-wise pruning, which halts the LLM forward pass when the next token certainty is high in the intermediate layers. This approach can be also thought as a way of reducing the effective depth of the model at test time. In contrast, quantization encompasses the transformation of floating-point numerical representations (predominantly FP32) into reduced-precision integer formats, such as INT8 or INT4 [21, 15, 22]. When implemented on hardware accelerators, these lower-precision representations facilitate superior memory bandwidth utilization, addressing a primary bottleneck in modern large-scale models [23]; moreover, integer-based computations yield enhanced processing speed and substantially improved energy efficiency [24]. Finally, parallelization techniques during inference, such as tensor and pipeline parallelism, enable the distribution of computational workload across multiple accelerators, thereby reducing latency and increasing throughput, although this often requires careful consideration of communication overhead and load balancing [25, 26].
Parallel attention-feedforward fusion. GPT-J[27] introduced a parallel formulation of the transformer decoder layer, executing attention and feedforward sub-blocks concurrently:
$y=x+\text{MHA}(\text{LayerNorm}(x))+\text{MLP}(\text{LayerNorm}(x))$
For models trained with tensor parallelism, this modification halves the number of required all-reduce operations. Additionally, it reduces memory bandwidth usage by eliminating one read and write operation of hidden states from HBM. The input projections for the attention and MLP sub-blocks can also be fused into a single kernel, further increasing arithmetic density. Consequently, training time is reduced by approximately 15% without observable performance degradation. PaLM[28] also adopted this formulation, noting that negative effects from deviating from the standard transformer self-attention diminish with increasing model size. This parallel formulation continues to be employed in more recent LLMs, including Gemini 1.5 Flash[29].
We differ from these approaches by focusing specifically on already-trained models with the standard transformer formulation[6, 30, 31], rather than training a model from scratch. By rearranging the computational graph, we merge two consecutive transformer-decoder layers by first executing both attention subblocks and both feedforward subblocks sequentially (Figure 2). Similar to GPT-J and PaLM, the primary source of acceleration in our method arises from halving the number of all-reduce operations across the layers where we apply LP.
@@
Parallelism via Computational Graph Optimization. Recent research has investigated architectural layer-level optimization strategies to enhance transformer model inference efficiency. The Staircase Transformer [32] implements parallel layer execution with dynamic recurrent computation based on model requirements. Similarly, the Staggering Transformer [33] achieves layer parallelization by connecting layer $l_{k}$ at time step $t$ to both the $(t-1)$ output of layer $l_{k-1}$ and the $t$ output of layer $l_{k-2}$. To the best of our knowledge, no research has addressed the fusion of consecutive layers through tensor parallelism.
@@ 3 Effective Depth
In this section we investigate the effective depth of pretrained LLMs by applying several transformations and measuring the resulting perplexity degradation. We reveal loose dependencies between intermediary layers. The transformations consist of shuffling, merging, and pruning transformer layers. To avoid the combinatorial explosion resulting from considering all the possible subsets of transformer layers, we instead apply our transformations to all the contiguous stretch of layers. If $L=\{\ell_{1},\ldots,\ell_{N}\}$ are the ordered layers, then we apply our transformations to all the sub-list $\{\ell_{i}\}_{i=s}^{e}$ with $1\leq s\leq e\leq N$. Previous works have shown that—at least when considering pruning—the importance of layers is well behaved, with low importance layers close to one another [34], which justifies considering contiguous stretch of layers only.
Shuffling blocks. We start by shuffling contiguous stretches of layers, re-ordering the layers according to random permutations. Results are shown in Fig. 3.(a). While shuffling the early and two last layers significantly raises the perplexity, there are large stretches of blocks which can be shuffled with surprisingly low impact on the perplexity. For instance, one can shuffle the layers $\{\ell_{i}\}_{15\leq i<25}$ of Llama 2 7B and only have a increase in perplexity of $2.9$. This goes against the classical belief that models build deeply hierarchical representations, where features in previous layers are leveraged to build more complex features in later layers. We interpret this as multiple layers working at the same level of abstraction. Using these insights, we define the effective depth of an LLM as the shortest depth required to efficiently leverage existing latent representations without a significant loss of performance. This reveals an important level of layer decoupling within the model.
Running blocks in parallel. The observed layer decoupling suggests that specific transformer operations may be executed independently, providing an opportunity for parallel computation. More precisely, let’s consider two sequential transformer layers $\ell_{k}$ and $\ell_{k+1}$, each comprising attention and Feed-Forward Network (FFN) sub-blocks ($A_{k}(\cdot)$ and $F_{k}(\cdot)$, respectively). The standard sequential output ${\bm{y}}$ for these layers, given an input ${\bm{x}}$, is given by:
@@
Contiguous 2-parallel. Instead of parallelizing long stretches of layers, we experiment with running pairs of consecutive layers in parallel. This springs from the assumption that local ordering matters less than global ordering, i.e. shuffling consecutive layers introduces fewer potential issues than shuffling layers separated by larger distances. As an example, if we apply the proposed transformation to layers $\{\ell_{15},\ell_{16},\ell_{17},\ell_{18},\ell_{19}\}$, it would result in the following process: (1) the two layers $\{\ell_{15},\ell_{16}\}$ process the input in parallel (according to equation (PAR)), (2) the output is forwarded to layers $\{\ell_{17},\ell_{18}\}$ which process it in parallel; finally, in (3) their joint output is fed to layer $\ell_{19}$ which processes it on its own as any normal layer. The effect of such transformation of the compute graph on the perplexity can be seen in Fig. 3(e). Remarkably, it is possible to run wide stretches of consecutive pairs of blocks in parallel with only a minor degradation of perplexity. For instance, one can apply this transformation from layer $4$ to layer $29$ with only a degradation of perplexity of $2.9$, while reducing the model depth from $32$ to $19$. The success of this approach led us to also try running triplets of consecutive layers in parallel, but we found it to perform worse.
Exploring other transformations. We also experiment with pruning (Fig. 3(b)) and merging (Fig. 3(c)). Pruning has already been studied in several prior works [35, 36]. While pruning a layer leads to a larger perplexity increase compared to parallelizing two blocks, it offers the advantage of reducing the model’s parameter count, which results in improved throughput and more efficient memory usage.